You are given the following information for Cleen Power Co. Assume the company’s tax rate is 40 percent. Debt: 10,000 7.7 percent coupon bonds outstanding, $1,000 par value, 15 years to maturity, selling for 106 percent of par; the bonds make semiannual payments. Common stock: 490,000 shares outstanding, selling for $67 per share; the beta is 1.10. Market: 8 percent market risk premium and 5.7 percent risk-free rate. What is the company's WACC? (Do not round intermediate calculations.

WACC % ?
Answer #1
A B C D E F G
Shares/Bonds Price Market Value Rate Weight Weight*Rate
Debt 10,000 =1000*106%=1060 =10000*1060=10600000 =RATE(15*2,((7.7%/2)*1000),-C2,1000)*2=7.05% =D2/SUM(D2:D3)=24.41% =(E2*F2)*(1-40%)=1.03%
Common Stock 490,000 67 =490,000*67=32830000 =5.7%+1.1*8%=14.50% ==D3/SUM(D2:D3)=75.59% =E3*F3=10.96%
11.99%
WACC=SUM(G2:G3)=11.99%